OpenStudy (anonymous):

The following list shows the items and prices for a restaurant order. Calculate the total amount if there is a 7% tax and the customer leaves a 20% gratuity. 1 appetizer: $7.39 2 entrees: $16.99 each 1 dessert: $5.29 2 drinks: $2.30 each a. $65.82 b. $65.10 c. $54.85 d. $40.60

Vocaloid (vocaloid):

let's leave the tax calculations aside for a minute, and just look at the cost of the items if he ordered 2 entries at 16.99 each, how much did he pay for the entries?

OpenStudy (anonymous):

33.98

Vocaloid (vocaloid):

great! he ordered 2 drinks at 2.30 each, how much did he pay for the drinks?

OpenStudy (anonymous):

4.6

Vocaloid (vocaloid):

great! so, he paid 4.60 for the drinks, 33.98 for the entries. he also bought 1 appetizer for 7.39 and 1 dessert for 5.29, how much did he pay for all these items?

OpenStudy (anonymous):

51.26

Vocaloid (vocaloid):

good so far! however, we're not finished yet he paid 51.26 for the items, but he also paid a 7% tax and a 20% gratuity the 7% tax is equal to (0.07)*(cost of the items) how much is the tax?

OpenStudy (anonymous):

3.60 i rounded

Vocaloid (vocaloid):

good! now, the 20% gratuity = (0.20)*(cost of the items) what is the gratuity?

OpenStudy (anonymous):

.71764

OpenStudy (anonymous):

or do you do the orginal cost then multiply

OpenStudy (anonymous):

10.30

Vocaloid (vocaloid):

oh, whoops, sorry, I forgot that gratuity gets calculated after tax... so, let's add the cost of the items + tax cost of the items = 51.26 tax = 3.60 let's add these together to get 54.86, with me so far?

OpenStudy (anonymous):

yea

Vocaloid (vocaloid):

alright, gratuity = 0.20*(54.86) = 10.972 gratuity + total cost = 10.972 + 54.86 = $65.82 (rounded)

Vocaloid (vocaloid):

that should be our final answer, sorry for the confusion > >

OpenStudy (anonymous):

lol its okay i have another i wont u too answer